OpenStudy (anonymous):

Line AB contains points A (1, 2) and B (−2, 6). The slope of line AB is

OpenStudy (mathmale):

Please look up the formula for the slope of a line connecting two points. I did an Internet search and came up with: https://www.google.com/search?q=slope+of+the+line+connecting+two+points&rlz=1C1CHFX_enUS461US461&oq=slope+of+the+line+connecting+two+points&aqs=chrome..69i57j0l5.7097j0j7&sourceid=chrome&espv=210&es_sm=122&ie=UTF-8

OpenStudy (anonymous):

Thanks but im still not sure what nswer to pick could you help me out?zero undefined positive negative

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ 1}}\quad ,&{\color{blue}{ 2}})\quad B&({\color{red}{ -2}}\quad ,&{\color{blue}{ 6}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (mathmale):

@SwazeyMark: You might not agree right now, but the best way in which anyone could help you right now is to help you learn to find slopes yourself using the formula given by jdoe0001.

OpenStudy (mathmale):

You will likely need this formula again and again.

OpenStudy (mathmale):

Hint: y2-y1=6-2=4. Your turn: x2-x1 = ??

OpenStudy (anonymous):

-2-1 right

OpenStudy (mathmale):

Great! then the slope you wanted is the ratio of 4 to -3: \[m= \frac{ -4 }{ 3 }\]

OpenStudy (mathmale):

And that's it. thanks for your patience and persistence.

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (mathmale):

My great pleasure, Mark! Hope to meet you again on OpenStudy!

OpenStudy (anonymous):

so it's negative right?